CHS is a large multidivision firm. One division, Health Services, is well known inside CHS for its efficient information technology (IT). A smaller division, Optics, has approached Health Services with a proposal that it provide IT support in the form of machine time for some of Optics’s billing and administrative work.

 

After an analysis of the demands that Optics would place on the system, the IT manager of Health Services notes that Health Services would have to lease a new server because of the additional load. The lease rates for the current server are a fixed annual lease of $4,200 and it averages machine time of 3,800 hours annually. The new server leases for an annual rate of $6,900. Because the new server is a faster machine, Health Services can complete its current requirements in only 2,450 hours. The work for Optics is estimated to be 2,000 hours.

 

In addition to leasing a new server, there are two other changes Health Services would have to make in IT. First, it will have to upgrade its server support position. The IT manager estimates that it will cost an additional $32,000 per year to get an individual with the necessary advanced training. In addition, Health Services has a contract for service from the machine vendor. The support contract is a fixed-price contract of $2 per hour of machine usage. The current lease contract can be canceled at no cost if Health Services leases a more expensive machine.

 

Required:

a. Assume that no outside market exists for this service and that Health Services would have excess capacity on the new server. What is the optimal transfer price rule Health Services should use to charge Optics?

b. Suppose Optics uses 2,000 hours on the new machine. What is the average cost per hour Optics would pay using the rule you developed in requirement (a)?

c. Suppose Optics uses 200 hours on the new machine. What is the average cost per hour Optics would pay using the rule you developed in requirement (a)?
Answer #1

4A) 6,900-4,200= 2,700 +32,000= 34,700

3,800-2450= 1,350* 2= 2,700

34,700- 2,700= 32,000 total fixed fee    variable cost= 2


4B) 2,000* 2= 4,000 

32,000+ 4,000= 36,000/ 2,000= 18 average cost per hour


4C) 200* 2= 400 

32,000+ 400= 32,400/ 200= 162 average cost per hour